OpenStudy (anonymous):

Given that the graph of f(x) passes through the point (6,10) and that the slope of its tangent line at (x,f(x)) is 2x+5, what is f(5)?

OpenStudy (amistre64):

what are your thoughts?

OpenStudy (anonymous):

not sure how to start it

OpenStudy (danjs):

apply amistre's profile picture to your derivative function f ' (x ) = 2x + 5, to return to f(x)

OpenStudy (danjs):

the given point is needed to figure what the integration constant is.

OpenStudy (danjs):

integrate f '(x) with respect to x \[\int\limits f'(x)*dx = \int\limits [2x + 5]*dx\]